# NeuroRVQ: Multi-Scale EEG Tokenization for Generative Large Brainwave Models

### v2

@@ Abstract
Electroencephalography (EEG) captures neural activity across multiple temporal and spectral scales, yielding signals that are rich but complex for representation learning. Recently, EEG foundation models trained to predict masked signal-tokens have shown promise for learning generalizable representations. However, their performance is hindered by their signal tokenization modules. Existing neural tokenizers fail to preserve high-frequency dynamics, limiting their ability to reconstruct EEG signals with high fidelity. We introduce NeuroRVQ, a scalable Large Brainwave Model (LBM) centered on a codebook-based tokenizer. Our tokenizer integrates: (i) multi-scale feature extraction modules that capture the full frequency neural spectrum; (ii) hierarchical residual vector quantization (RVQ) codebooks for high-resolution encoding; and, (iii) an EEG signal phase- and amplitude-aware loss function for efficient training. This design enables efficient EEG compression while supporting accurate reconstruction across all frequency bands, leading to robust generative masked modeling. Our empirical results demonstrate that NeuroRVQ achieves lower reconstruction error and outperforms existing LBMs on a variety of downstream tasks. More broadly, the NeuroRVQ tokenizer establishes a strong prior for codebook-based general-purpose brainwave models, enabling advances in neural decoding, generative modeling and multimodal biosignal integration.
@@ 1 Introduction
Brain activity unfolds across multiple frequency scales. Brainwave signals can be categorized into distinct frequency bands—namely, delta (0.5–4 Hz), theta (4–8 Hz), alpha (8–13 Hz), beta (13–30 Hz) and gamma (>30 Hz)—each linked to specific cognitive and physiological states. For example, delta waves dominate during deep sleep, while alpha waves emerge during relaxing states (especially with closed eyes) but tend to decrease in amplitude during movement. Together, these bands provide a dynamic window into the brain’s activity and internal states. EEG analysis must capture the multiscale structure of brainwaves, both slow and fast dynamics, to fully represent neural information.
A range of techniques, from advanced signal processing to machine learning, have been employed for brainwave decoding. More recently, Large Brainwave Models (LBMs) offer strong generalization capabilities without relying on BCI task-specific data collection and model training. These models are typically trained using self-supervised masked learning: EEG signals are discretized into tokens; and, during training, tokens are intentionally masked so that models learn to reconstruct them. However, reconstructing raw EEG signals remains a significant challenge due to their inherently noisy and complex nature. Tokenizers must effectively capture spatio-temporal patterns (e.g., location of recording electrodes, time and frequency) to convert them into meaningful token representations. Current tokenizers struggle to preserve the complete structural information (especially high frequency components) necessary for robust generative masked modeling. The key to unlocking foundation-scale masked modeling for EEG lies in the tokenizer:
A well-designed tokenizer should not only compress continuous neural signals into discrete tokens but also enable faithful reconstruction of the original waveform across all important frequency scales.
@@
    captures multi-scale frequency features by applying temporal convolutions with varying kernel sizes;
@@
    encodes hierarchical representations in Residual Vector Quantization (RVQ) codebooks, one per frequency scale, enabling us to capture the complex patterns necessary for high-fidelity signal reconstruction in 32 codebooks ($2^{32}$ parameters);
@@
    employs a training procedure grounded in strong signal processing principles that captures both amplitude and wrapped phase information of the EEG signal via sine and cosine representations.
We evaluate the efficacy of our tokenizer by comparing NeuroRVQ with state-of-the-art LBMs. NeuroRVQ achieves up to 15% higher accuracy on four BCI classification tasks, demonstrating the effectiveness of codebook-based modeling when codebooks faithfully reconstruct brain signals.
@@ 2 Background & Related Work
The roots of EEG analysis can be traced back to classical neuroscience methods, when handcrafted features such as Power Spectral Density (PSD) bands and Independent Component Analysis (ICA) were regarded as standard (Bashashati et al., 2007; Handy, 2009; Rao, 2013; Nam et al., 2018; McFarland et al., 2006). While these methods provided valuable insights into neural dynamics, they often failed to generalize due to high inter-subject variability, the noisy nature of EEG signal and their limited capacity to capture the full complexity of brain activity.
The deep learning era marked a shift towards data-driven feature extraction: brainwave decoders such as EEGNet (Lawhern et al., 2018), EEGInception (Santamaría-Vázquez et al., 2020), EEGConformer (Song et al., 2023), and Brainware-Scattering Net (Barmpas et al., 2023) reduced reliance on handcrafted features by automatically learning representations of brain activity. To address domain-specific challenges such as subject variability, researchers incorporated domain-specific knowledge into these architectures, such as subject-alignment to adjust feature representations (Bakas et al., 2022; Wei et al., 2022; Barmpas et al., 2024b). Despite these advances, most deep learning models remain limited: they require carefully labeled data and extensive task-specific training, hindering their ability to scale and generalize across diverse BCI tasks.
@@
Tokenization is critical for robust generative masked modeling, yet current tokenizers struggle to capture the full structure of brain signals. NeuroGPT relies on a deep backbone-based tokenizer while other models like BIOT and EEGPT on transformer-based tokenizers, which often learn entangled and redundant representations not optimized for reconstruction. As a result, the generated tokens lack the fidelity required for masked prediction. In contrast, LaBraM adopts a codebook-based approach to learn reconstructable neural tokens. However, while discrete codebook-based tokenizers such as VQ-VAE (Esser et al., 2020) work well in computer vision, directly adopting them for brain signals have not yielded faithful signal reconstruction—until NeuroRVQ.
@@ 3.1 Model Architecture
NeuroRVQ tokenizer is a specialized network designed to convert raw EEG signals into a sequence of compact and informative neural tokens. This transformation reduces the inherently high-dimensional and noisy nature of EEG into a structured lower-dimensional representation that preserves essential temporal–spectral patterns. In doing so, the tokenizer provides a kind of "neural grammar" for brain activity.
Figure 1 illustrates the tokenizer architecture and processing flow. The input multi-variate timeseries is first segmented into EEG patches. These patches are encoded by the multi-scale temporal encoder, that captures features in multiple resolutions and are then combined via the transfromer encoder. For each scale, RVQ codebooks discretize the embeddings into a sequence of neural tokens. Finally, these tokens are combined and passed through the tokenizer decoder to reconstruct the input EEG patches using the Fourier spectrum.
Generating Patches: Let $X\in\mathbb{R}^{C\times T}$ denote the input EEG signal, where $T$ is the number of time points and $C$ is the number of electrodes. The signal is first segmented into temporal patches. To ensure that the tokenizer can accommodate EEG signals with arbitrary numbers of channels and variable durations, we adopt the following approach: during model pre-training, each input multi-variate EEG signal is represented by $P$ patches of length $w$ (corresponding to a 1-second time window). This results in a segmented input sample $\mathbf{x}\in\mathbb{R}^{P\times w}$ which provides flexibility to handle heterogeneous EEG recordings and facilitates robust pre-training across diverse datasets.
@@
RVQ Codebook: To efficiently discretize the multi-scale patch representations $\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{S}$ while preserving fine-grained information of the EEG signal, we use $S$ Residual Vector Quantization (RVQ) codebooks. For each temporal branch, the RVQ codebook $\mathcal{R}$ is defined as:
@@
$\mathbf{z_{1}}=\arg\min_{\mathbf{v}\in\mathcal{V}_{1}}\left\|l_{2}(\mathbf{p}^{1})-l_{2}(\mathbf{v})\right\|$ (3)
@@
Tokenizer Decoder: The purpose of the decoder module (a series shared transformer layers followed by prediction heads) is to reconstruct the original signal information based on the learned codebook tokens. Directly reconstructing raw EEG signals has been shown to yield unstable training and poor performance, due to the inherently noisy and high-dimensional nature of EEG. Following the strategy of LaBraM (Jiang et al., 2024), we instead leverage the Fourier spectrum (more specifically the phase $\phi$ and spectral amplitude $A$) as a more structured target for reconstruction.
@@ 3.2 Tokenizer Training
Naively treating the sine and cosine components independently by applying mean squared error to each dimension separately during training fails to explicitly capture the circular nature of phase: minimizing the MSE of sine and cosine individually does not guarantee that the resulting vector lies on the unit circle and it may allow predictions that deviate from valid phase angles. To address these limitations, we introduce a unit-circle-aware phase loss $\mathcal{L}_{\text{unit-loss}}$ that directly optimizes the directional alignment of the predicted phase vector with the target phase while enforcing unit-norm constraints. Specifically, we represent each phase as a 2D vector ($\cos\phi,\sin\phi$) and compute the cosine similarity between predicted and target vectors, combined with a small regularization term that penalizes deviations from unit length:
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{L}_{\text{unit-loss}}}=1-\underbrace{\sum_{i}\frac{\cos{\hat{\phi_{i}}}\cos{\phi_{i}}+\sin{\hat{\phi_{i}}}\sin{\phi_{i}}}{\sqrt{\cos{\hat{\phi_{i}}}^{2}+\sin{\hat{\phi_{i}}}^{2}}\,\sqrt{\cos{\hat{\phi_{i}}}^{2}+\sin{\hat{\phi_{i}}}^{2}}}}_{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\text{Cosine Similarity}}}\;+\;\underbrace{\lambda_{circle}\cdot\sum_{i}\big(\cos{\hat{\phi_{i}}}^{2}+\sin{\hat{\phi_{i}}}^{2}-1\big)^{2}}_{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\text{Unit-Circle Penalization Term}}}$ (6)
(a) Prediction not constrained to unit circle.
(b) Prediction constrained to $|v|=1$.
This formulation has several advantages: it naturally respects the circular topology of phase, avoids discontinuities near angular boundaries and ensures that all predicted vectors correspond to valid phase angles. The phases are essentially represented as unit vectors in 2D space, making the errors between them more geometrically meaningful.
@@
Finally, although reconstructing raw EEG waveforms is unstable, we found that incorporating a temporal-domain reconstruction loss provides additional guidance. Specifically, we transform the predicted log-amplitude and phase back to the time domain (the predicted log-amplitudes are exponentiated and rescaled to recover the original amplitude spectrum before inverse Fourier transformation) and compute a temporal MSE term. The NeuroRVQ tokenizer is thus trained end-to-end with the following objective:
@@
$\mathcal{L}_{T}=\sum_{i}\underbrace{\left\|\log(1+\hat{A_{i}})-\log(1+A_{i})\right\|_{2}^{2}}_{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\text{Log-Amplitude Loss}}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{L}_{\text{unit-loss}}}+\underbrace{\left\|\hat{X_{i}}-X_{i}\right\|_{2}^{2}}_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\text{Temporal Loss}}}+\mathcal{L}_{Q},$ (7)
@@
where $\hat{A}$, $\sin\hat{\phi}$, and $\cos\hat{\phi}$ denote reconstructed Fourier components, $\hat{X}$ denotes reconstructed EEG in the time domain and $\mathcal{L}_{Q}$ is the quantization loss as defined in Esser et al. (2020).
@@ 4 NeuroRVQ Foundation Model
Figure 3 shows the architecture of NeuroRVQ foundation model. The architecture consists of a multi-scale temporal encoder (Section 3.1) that extracts hierarchical temporal features across multiple resolutions, followed by a series of transformer layers that model long-range complex dependencies. A series of token prediction heads reconstructs the masked tokens from the contextualized embeddings using the trained $\mathcal{R}$ codebooks, which are used only during the pre-training of the backbone model. In line with LaBraM, a symmetric masking strategy is employed, generating the spatially inverse of each original mask during training to enhance efficiency and data diversity. By combining hierarchical temporal encoding, transformer-based context modeling and symmetric masking, the foundation model effectively captures rich spatio-temporal patterns and produces high-fidelity token predictions. For downstream tasks, the pre-trained backbone is extended with task-specific classifications heads.

### v3

@@ Abstract
Electroencephalography (EEG) captures neural activity across multiple temporal and spectral scales, yielding signals that are rich but complex for representation learning. EEG foundation models trained to predict masked signal tokens have shown promise in learning generalizable representations. However, their performance is hindered by signal tokenization: existing EEG signal tokenizers fail to preserve high-frequency dynamics, limiting their ability to reconstruct signals with high fidelity.
We introduce NeuroRVQ, a scalable Large Brainwave Model (LBM) centered on a codebook-based tokenizer. Our tokenizer integrates: (i) multi-scale feature extraction modules that capture the full frequency neural spectrum; (ii) hierarchical residual vector quantization (RVQ) codebooks for high-resolution encoding; and (iii) a unit circle phase-aware loss for efficient training. This design enables efficient EEG compression while supporting accurate reconstruction across all frequency bands, leading to robust generative masked modeling. Our results demonstrate that NeuroRVQ achieves a lower reconstruction error and outperforms existing LBMs on a variety of downstream tasks. Preliminary results on electromyography (EMG) and electrocardiogram (ECG) signal reconstruction suggest NeuroRVQ’s potential to extend beyond EEG. More broadly, the NeuroRVQ tokenizer establishes a strong prior for codebook-based general-purpose brainwave models, enabling advances in neural decoding, generative modeling and multimodal biosignal integration.
Machine Learning, ICML, EEG, BCIs
@@ 1 Introduction
Brain activity unfolds across multiple frequency scales. Brainwave signals can be categorized into distinct frequency bands (Newson and Thiagarajan, 2018)—namely, delta (0.5–4 Hz), theta (4–8 Hz), alpha (8–13 Hz), beta (13–30 Hz) and gamma (>> 30 Hz)—each linked to specific cognitive and physiological states. For example, delta waves dominate during deep sleep, while alpha waves emerge during relaxing states (especially with closed eyes), but tend to decrease in amplitude during movement. Together, these bands provide a dynamic window into the brain’s activity and internal states. EEG analysis must capture the multiscale structure of brainwaves, both slow and fast dynamics, to fully represent neural information.
A range of techniques, from advanced signal processing to machine learning, have been employed for brainwave decoding. More recently, Large Brainwave Models (LBMs) offer strong generalization capabilities without relying on BCI task-specific data collection and model training. These models are typically trained using self-supervised masked learning: EEG signals are discretized into tokens; and, during training, tokens are intentionally masked so that models learn to reconstruct them. However, reconstructing raw EEG signals remains a significant challenge due to their inherently noisy and complex nature. Tokenizers must effectively capture spatio-temporal patterns (e.g., location of recording electrodes, time and frequency) to convert them into meaningful token representations. Current tokenizers struggle to preserve the complete structural information (especially high frequency components) necessary for robust generative masked modeling.
The key to unlocking foundation-scale masked modeling for EEG lies in the tokenizer:
A well-designed EEG tokenizer should not only compress continuous neural signals into discrete tokens but also enable faithful reconstruction of the original waveform across all important frequency scales.
@@
    captures multi-scale frequency features by applying temporal convolutions with varying kernel sizes, to effectively disentangle the underlying frequency components;
@@
    encodes hierarchical representations in Residual Vector Quantization (RVQ) codebooks, one per frequency scale, enabling us to learn and preserve the complex patterns necessary for high-fidelity signal reconstruction;
@@
    introduces a unit circle phase-aware training loss, grounded in strong signal processing principles, that captures both amplitude and wrapped phase information of the EEG signal via sine and cosine representations; and
  4. (4)
    can be adapted to other biosignal modalities by tuning kernel sizes for modality-specific frequencies and scaling RVQ codebooks to signal complexity.
We evaluate the efficacy of the NeuroRVQ tokenizer by comparing its reconstruction capabilities to other state-of-the-art tokenizers. NeuroRVQ foundation model achieves up to 15% higher performance on five BCI downstream tasks compared to other EEG foundation models, demonstrating the effectiveness of codebook-based modeling when codebooks faithfully reconstruct brain signals.
@@ 2 Background & Related Work
The roots of EEG analysis can be traced back to classical neuroscience methods, when handcrafted features such as Power Spectral Density (PSD) bands and Independent Component Analysis (ICA) were regarded standard (Bashashati et al., 2007; Handy, 2009; Rao, 2013; Nam et al., 2018; McFarland et al., 2006). Although these methods provided valuable insights into neural dynamics, they often failed to generalize due to high inter-subject variability, the noisy nature of EEG signal, and their limited capacity to capture the full complexity of brain activity.
The deep learning era marked a shift towards data-driven feature extraction: brainwave decoders such as EEGNet (Lawhern et al., 2018), EEGInception (Santamaría-Vázquez et al., 2020), EEGConformer (Song et al., 2023) and Brainware-Scattering Net (Barmpas et al., 2023) reduced reliance on handcrafted features by automatically learning representations of brain activity. To address domain-specific challenges such as subject variability, researchers incorporated domain-specific knowledge into these architectures, such as subject-alignment to adjust feature representations (Bakas et al., 2022; Wei et al., 2022; Barmpas et al., 2024b). Despite these advances, most deep learning models remain limited: they require carefully labeled data and extensive task-specific training, hindering their ability to scale and generalize across diverse BCI tasks.
@@
Tokenization is critical for robust generative masked modeling, yet current tokenizers struggle to capture the full structure of brain signals. NeuroGPT relies on a deep backbone-based tokenizer while other models like BIOT and EEGPT on transformer-based tokenizers, which often learn entangled and redundant representations not optimized for reconstruction. As a result, the generated tokens lack the fidelity required for masked prediction. In contrast, LaBraM and more recently BrainOmni  (Xiao et al., 2025) adopt a codebook-based approach to learn reconstructable neural tokens. However, while discrete codebook-based tokenizers work well in computer vision and audio, directly adopting them for brain signals has not yielded faithful signal reconstruction—until NeuroRVQ.
@@ 3.1 Model Architecture
NeuroRVQ tokenizer is a specialized network designed to convert raw EEG signals into a sequence of compact and informative neural tokens. This transformation reduces the inherently high-dimensional and noisy nature of EEG into a structured lower-dimensional representation preserving essential temporal–spectral patterns. In doing so, the tokenizer provides a kind of ”neural grammar” for brain activity.
Figure 1 illustrates the tokenizer architecture and processing flow. The input multi-variate timeseries is first segmented into EEG patches. These patches are encoded by the multi-scale temporal encoder, that captures features in multiple resolutions and are then combined via the transfromer encoder. For each scale, RVQ codebooks discretize the embeddings into a sequence of neural tokens. Finally, these tokens from all scales are combined and passed through the tokenizer decoder to reconstruct the input EEG patches using the Fourier spectrum.
Generating Patches:
Let $X\in\mathbb{R}^{C\times T}$ denote the input EEG signal, where $T$ is the number of time points and $C$ is the number of electrodes. The signal is first segmented into temporal patches. To ensure that the tokenizer can accommodate EEG signals with arbitrary numbers of channels and variable durations, we adopt the following approach: during model pre-training, each input multi-variate EEG signal is represented by $P$ patches of length $w$ (corresponding to a 1-second time window). This results in a segmented input sample $\mathbf{x}\in\mathbb{R}^{P\times w}$ which provides flexibility to handle heterogeneous EEG recordings and facilitates robust pre-training across diverse datasets.
@@
(a) Prediction not constrained to unit circle.
(b) Prediction constrained to unit circle.
RVQ Codebook: To efficiently discretize the multi-scale patch representations $\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{S}$ while preserving fine-grained information of the EEG signal, we use $S$ Residual Vector Quantization (RVQ) codebooks (Gray, 1984). For each temporal branch, the RVQ codebook $\mathcal{R}$ is defined as:
@@
$\mathbf{z_{1}}=\arg\min_{\mathbf{v^{1}}\in\mathcal{V}_{1}}\left\|l_{2}(\mathbf{p}^{1})-l_{2}(\mathbf{v^{1}})\right\|$ (3)
@@
Tokenizer Decoder: The purpose of the decoder module (a series shared transformer layers followed by prediction heads) is to reconstruct the original signal information based on the learned codebook tokens. Directly reconstructing raw EEG signals has been shown to yield unstable training and poor performance, due to the inherently noisy and high-dimensional nature of EEG. Following the strategy of LaBraM (Jiang et al., 2024), we leverage the Fourier spectrum (more specifically the phase $\phi$ and spectral amplitude $A$) as a more structured target for reconstruction.
@@ 3.2 Tokenizer Training
Naively treating the sine and cosine components independently by applying mean squared error to each dimension separately during training fails to explicitly capture the circular nature of phase: minimizing the MSE of sine and cosine individually does not guarantee that the resulting vector lies on the unit circle and it may allow predictions that deviate from valid phase angles. To address these limitations, we introduce a unit circle phase-aware loss $\mathcal{L}_{\text{unit-loss}}$ that directly optimizes the directional alignment of the predicted phase vector with the target phase while enforcing unit-norm constraints. Specifically, we represent each phase as a 2D vector ($\cos\phi,\sin\phi$) and compute the cosine similarity between the predicted and target vectors, combined with a small regularization term that penalizes deviations from unit length:
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{L}_{\text{unit-loss}}}$ $=1-\underbrace{\sum_{i}\frac{\cos{\hat{\phi_{i}}}\cos{\phi_{i}}+\sin{\hat{\phi_{i}}}\sin{\phi_{i}}}{\sqrt{\cos^{2}{\hat{\phi_{i}}}+\sin^{2}{\hat{\phi_{i}}}}\,\sqrt{\cos^{2}{\phi_{i}}+\sin^{2}{\phi_{i}}}}}_{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\text{Cosine Similarity}}}$ (6)
$\quad+\;\underbrace{\lambda_{circle}\cdot\sum_{i}\big(\cos^{2}{\hat{\phi_{i}}}+\sin^{2}{\hat{\phi_{i}}}-1\big)^{2}}_{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\text{Unit-Circle Penalization Term}}}$
@@
This formulation has several advantages: it naturally respects the circular topology of phase, avoids discontinuities near angular boundaries and ensures that all predicted vectors correspond to valid phase angles. The phases are essentially represented as unit vectors in 2D space, making the errors between them more geometrically meaningful. Finally, although reconstructing raw EEG waveforms is unstable, we found that incorporating a temporal-domain reconstruction loss provides additional guidance. Specifically, we transform the predicted log-amplitude and phase back to the time domain (the predicted log-amplitudes are exponentiated and rescaled to recover the original amplitude spectrum before the inverse Fourier transformation) and compute a temporal MSE regularization term. The NeuroRVQ tokenizer is thus trained end-to-end with the objective:
@@
$\mathcal{L}_{T}$ $=\sum_{i}\underbrace{\left\|\log(1+\hat{A_{i}})-\log(1+A_{i})\right\|_{2}^{2}}_{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\text{Log-Amplitude Loss}}}$ (7)
$\quad+\;\underbrace{\left\|\hat{X_{i}}-X_{i}\right\|_{2}^{2}}_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\text{Temporal Loss}}}+\;{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{L}_{\text{unit-loss}}}\;+\;\mathcal{L}_{Q}$
@@
where $\hat{A}$, $\sin\hat{\phi}$, and $\cos\hat{\phi}$ denote reconstructed Fourier components, $\hat{X}$ denotes reconstructed EEG in the time domain and $\mathcal{L}_{Q}$ is the commitment loss.
@@ 4 NeuroRVQ Foundation Model
Figure 3 shows the architecture of NeuroRVQ foundation model. The architecture consists of a multi-scale temporal encoder (Section 3.1) that extracts hierarchical temporal features across multiple resolutions, followed by a series of transformer layers that model long-range complex dependencies. A series of token prediction heads reconstructs the masked tokens from the contextualized embeddings using the trained $\mathcal{R}$ codebooks, which are used only during the pre-training of the backbone model. In line with LaBraM (Jiang et al., 2024), a symmetric masking strategy is employed, generating the spatially inverse of each original mask during training to enhance efficiency and data diversity. By combining hierarchical temporal encoding, transformer-based context modeling and symmetric masking, the foundation model effectively captures rich spatio-temporal patterns and produces high-fidelity token predictions. For downstream tasks, the pre-trained backbone is extended with task-specific classifications heads.